Find the measure of an angle such that the difference between its supplement and 2 times its complement is 37
Answer by macston(5194)

x=measure of angle
(180 degrees-x)-2(90 degrees-x)=37 degrees
(180 degrees-x-180 degrees+2x=37 degrees
x=37 degrees ANSWER: The measure of the angle is 37 degrees.
CHECK:
(180 degrees-37 degrees)-2(90 degrees-37 degrees)=37 degrees
143 degrees-2(53 degrees)=37 degrees
143 degrees-106 degrees=37 degrees
37 degrees=37 degrees